1. at 20°C, The rate of a particular reaction is 0.001 M/s; what is the rate expected to be at 60°C, assuming all other factors remain the same?

2. Two compartments, labeled A and B, are separated by semipermeable membrane. What happens if the compartment A is filled with honey (a solution of sugar dissolved in water) and apartment B is filled with distilled water?

3. Which statement best explains why the following reaction is not likely to occur in a single step? 2C6H6(1)+15 O2(g)---> 12 CO2(g) + 6 H2O(g)
A. Exothermic reactions typically require several steps.
B. Gases do not usually react directly with liquids or solids.
C. The 17-body collision required is extremely unlikely.
D. The reaction is spontaneous and is not reversible.
1
Expert's answer
2017-10-26T15:26:06-0400
For many reactions happening around the room temperature, the rate of reaction doubles for every 10°C rise in temperature. This general rule can be expressed mathematically as follows:
k_T=k_(T^' )·2^(((T-T^')/10) )
Thus,
k_(60°C)=0.001(M s^(-1) )·2^(((60-20)/10))=0.016(M s^(-1) )
When the concentration of the solute in two compartments separated by semipermeable membrane is different, osmosis occurs. This means that the fluid (water in our case) will flow to the compartment where the concentration is lower. This flow will be proportional to the difference of concentration of sugar in honey and water in chamber B.
C. The 17-body collision required is extremely unlikely.